# TELEClass: Taxonomy Enrichment and LLM-Enhanced Hierarchical Text Classification with Minimal Supervision

### v2

@@ Acknowledgments
Research was supported in part by US DARPA INCAS Program No. HR001121C0165 and KAIROS Program No. FA8750-19-2-1004, National Science Foundation IIS-19-56151, and the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.
@@ 12             $T(c,c_{p})\leftarrow$ top terms ranked by affinity;
14      $T_{c}\leftarrow$ aggregate corpus-based and LLM-generated terms;
@@ 27 $\mathcal{D}^{\text{gen}}\leftarrow$ generate $q$ documents for each path using $\mathcal{G}$;
28 $F\leftarrow$ train classifier with $\mathcal{D}^{\text{core}}$ and $\mathcal{D}^{\text{gen}}$;
@@ 29 Return $F$;
Appendix B Experiment Setup
B.1 Compared Methods
  • Hier-0Shot-TC Yin et al. (2019) is a zero-shot approach, which utilizes a pretrained textual entailment model to iterative find the most similar class at each level for a document.
  • Hier-doc2vec Le and Mikolov (2014) is a weakly-supervised approach, which first trains document and class representations in a same embedding space, and then iteratively selects the most similar class at each level.
  • WeSHClass Meng et al. (2019) is a weakly-supervised approach using a set of keywords for each class. It first generates pseudo documents to pretrain text classifiers and then performs self-training.
  • TaxoClass Shen et al. (2021) is a weakly-supervised approach that only uses the class name of each class. It first uses a textual entailment model with a top-down search and corpus-level comparison to select core classes, which are then used as pseudo training data. We include both its full model, TaxoClass, and its variation TaxoClass-NoST that does not apply self-training on the trained classifier, which is the same as TELEClass.
B.2 Evaluation Metrics
@@
  • Example-F1 Sørensen (1948), which is also called micro-Dice coefficient, evaluates the multi-label classification results without ranking,
@@
    $\text{MRR}=\frac{1}{|\mathcal{D}|}\sum_{d_{i}\in\mathcal{D}}\frac{1}{|\mathbb{C}_{i}^{\text{true}}|}\sum_{c_{j}\in\mathbb{C}_{i}^{\text{true}}}\frac{1}{\min\{k|c_{j}\in\mathbb{C}^{\text{pred}}_{i,k}\}}.$ (5)
B.3 Implementation Details
We use Sentence Transformer Reimers and Gurevych (2019) all-mpnet-base-v2 as the text encoder for the similarity measure in Section 3.1 and Section 3.3. We query GPT-3.5-turbo-0125 for LLM-based taxonomy enrichment, core class annotation, and path-based generation. For corpus-based taxonomy enrichment, we get the term and class name embeddings using a pre-trained BERT-base-uncased Devlin et al. (2019), we select top $k=20$ enriched terms for each class (c.f. Section 3.2). We generate $q=5$ documents with path-based generation for each class. The document encoder in the final classifier is initialized with BERT-base-uncased for a fair comparison with the baselines. We train the classifier using AdamW optimizer with a learning rate 5e-5, and the batch size is 64. The experiments are run on one NVIDIA RTX A6000 GPU.
Appendix C Prompts for LLM
  • LLM-based enrichment for Amazon-531
    Instruction: [Target Class] is a product class in Amazon and is the subclass of [Parent Class]. Please generate 10 additional key terms about the [Target Class] that is relevant to [Target Class] but irrelevant to [Sibling Classes]. Please split the additional key terms using commas.
  • LLM-based enrichment for DBPedia-298
    Instruction: [Target Class] is a article category of Wikipedia articles and is the subclass of [Parent Class]. Please generate 10 additional key terms about the [Target Class] that is relevant to [Target Class] but irrelevant to [Sibling Classes]. Please split the additional key terms using commas.
  • Core class annotation for Amazon-531
  • Path-based generation for Amazon-531
  • Core class annotation for DBPedia-298
  • Path-based generation for DBPedia-298

### v3

@@ Acknowledgements.
Research was supported in part by US DARPA INCAS Program No. HR0011-21-C0165 and BRIES Program No. HR0011-24-3-0325, National Science Foundation IIS-19-56151, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.
@@ 12             $T(c,c_{p})\leftarrow$ top terms ranked by affinity;
14      $T_{c}\leftarrow$ aggregate corpus-based and LLM-generated terms Eq. 4;
@@ 27 $\mathcal{D}^{\text{gen}}\leftarrow$ generate $q$ documents for each path using $\mathcal{G}$;
28 $F\leftarrow$ train classifier with $\mathcal{D}^{\text{core}}$ and $\mathcal{D}^{\text{gen}}$ Eq. 6;
@@ 29 Return $F$;
Algorithm 1 summarizes the TELEClass method.
Appendix B Prompts for LLM
  • LLM-based enrichment for Amazon-531
    Instruction: [Target Class] is a product class in Amazon and is the subclass of [Parent Class]. Please generate 10 additional key terms about the [Target Class] that are relevant to [Target Class] but irrelevant to [Sibling Classes]. Please split the additional key terms using commas.
  • LLM-based enrichment for DBPedia-298
    Instruction: [Target Class] is an article category of Wikipedia articles and is the subclass of [Parent Class]. Please generate 10 additional key terms about the [Target Class] that are relevant to [Target Class] but irrelevant to [Sibling Classes]. Please split the additional key terms using commas.
  • Core class annotation for Amazon-531
  • Core class annotation for DBPedia-298
  • Path-based generation for Amazon-531
  • Path-based generation for DBPedia-298
Appendix C Evaluation Metrics
@@
  • Example-F1 (Sørensen, 1948), which is also called micro-Dice coefficient, evaluates the multi-label classification results without ranking,
@@
    (9) $\text{MRR}=\frac{1}{|\mathcal{D}|}\sum_{d_{i}\in\mathcal{D}}\frac{1}{|\mathbb{C}_{i}^{\text{true}}|}\sum_{c_{j}\in\mathbb{C}_{i}^{\text{true}}}\frac{1}{\min\{k|c_{j}\in\mathbb{C}^{\text{pred}}_{i,k}\}}.$